1) A stock has generated an annual average return of 9.5% with a standard deviation of 40.7% during the last 10 years. If the average risk-free rate was 1.7%, what was this stock's Sharpe Ratio? Round to two decimal places.

2) The standard deviation of a stock's annual returns is 40.4%. The standard deviation of market returns is 24.3%. If the correlation between the returns of the stock and the market is 0.3, what is this stock's beta? Round to two decimal places.
Answer #1

Answer :

1). Sharp Ratio = ( Rp - Rf ) / SDp

Where, Rp = Return on portfolio = 9.5%

Rf = Risk free return = 1.7%

SDp = Standard deviation of portfolio = 40.7%

Sharp Ratio = ( 9.5 - 1.7 ) / 40.7

= 7.8 / 40.7

= 0.19

2). beta of stock = correlation coefficient * standard deviation of stock / standard deviation of market

= 0.3 * ( 0.404 / 0.243 )

= 0.3 * 1.66

= 0.50